# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ Abstract
Method: We selected 146 relevant publications from the DBLP Computer Science Bibliography up to the end of 2024, following established systematic literature review guidelines. We explored publication trends, categorized them based on software engineering topics, and reported findings based on those topics.
Results: The most popular venues for publishing software engineering research on Jupyter notebooks are related to human-computer interaction instead of traditional software engineering venues. Researchers have addressed a wide range of software engineering topics on notebooks, such as code reuse, readability, and execution environment. Although reusability is one of the research topics for Jupyter notebooks, only 64 of the 146 studies can be reused based on their provided URLs. Additionally, most replication packages are not hosted on permanent repositories for long-term availability and adherence to open science principles.
@@ 1 Introduction
In recent years, Jupyter Notebook has emerged as a powerful and versatile computing platform for data science, scientific research, and software development [112, 135]. Its user-friendly interface has revolutionized the way researchers, developers, and data scientists conduct and communicate their work. Jupyter Notebook provides a unique platform for integrating code, visualizations, and narrative text, fostering reproducibility, and facilitating collaborative exploration of data and algorithms [171]. Seamlessly incorporating code, data, and output into a single file makes the Jupyter Notebook ideal for data analysis, scientific computing, and machine learning (ML) tasks. Moreover, its support for multiple programming languages, including Python, R, and Julia, makes it versatile and widely accessible to a diverse community of researchers, data scientists, and educators. However, despite their popularity, notebooks have been associated with several challenges, such as low reproducibility rates, problems with their execution environments, and excessive code duplication [46].
As Jupyter notebooks are often created and maintained by users without a software engineering background [150], there has been a growing interest from software engineering researchers to help notebook users integrate best software engineering practices into their notebooks. As a result, there has been a growing body of software engineering research that targets Jupyter Notebook. In this paper, we conduct a systematic literature review (SLR) of such research and the problems addressed by it.
We systematically searched academic publications to identify relevant studies related to software engineering research on Jupyter Notebook. We followed Kitchenham’s guidelines [71] for our SLR to ensure rigor and impartiality. We comprehensively searched all the papers on Jupyter Notebook indexed in the DBLP database [80] and published until 2024. Then, we filtered the papers focused on software engineering research on Jupyter notebooks. The review process involved several steps, including screening the titles and abstracts of the identified studies, assessing their relevance based on the inclusion criteria, and extracting relevant data from the selected studies. We finally selected 146 papers as our primary studies. During our SLR, we focused on the following two research questions (RQs):
@@
    RQ1: How much software engineering research on Jupyter Notebook has been published? We analyzed current trends in the publication of software engineering research on notebooks to help future researchers identify potential venues for their work. The number of publications has gradually increased over the years. Most (78 of 146 studies) software engineering research on notebooks has been published at conferences, indicating a fast-moving field. Our research found that other venues beyond core software engineering conferences published the highest number of studies; for example, the ACM Conference on Human Factors in Computing Systems (CHI) published 15 studies. Conversely, the International Conference on Software Engineering (ICSE) published six studies on software engineering research on notebooks. According to our study, a significant number of the articles we examined, specifically 31 out of 146 (21%), had affiliations with industry institutions such as Microsoft Research (12 studies) or IBM Research Lab (6 studies). This highlights the practical importance and real-world applications of the research findings. We identified 67 studies that provided URLs to replication packages. Of these, 64 URLs were deemed reusable, while three were not, due to missing source code or unclear execution instructions. Most replication packages (54 studies) are hosted in GitHub repositories, which is against the best practices for open science in software engineering [98] as such repositories can disappear over time. Only 13 replication packages were hosted in a permanent repository, such as Zenodo or Figshare.
@@
    RQ2: Which software engineering topics are being studied in software engineering research on Jupyter Notebook? We categorized the primary studies into 11 software engineering topics: code reuse and provenance, managing the computational environment and workflow, readability of notebooks, datasets of notebooks, documentation of notebooks, testing and debugging, visualization in notebooks, best practices, cell execution order, notebook code generation, and supporting other programming paradigms. Our findings indicate that code reuse and provenance is the most extensively researched topic related to Jupyter Notebook, with 32 studies focusing on it. This is followed by managing computational environment and workflow with 29 studies, and readability of notebooks with 11 studies. These results suggest that researchers are primarily concerned with code cells in notebooks, which often require significant human effort to understand. We also identified 19 publicly available Jupyter Notebook datasets used in the studies, most of which were sourced from Kaggle and GitHub repositories.
@@ 2.1.1 Cell-Based Structure
Code Cell: Code cells are the fundamental components of a Jupyter Notebook, where the source code (Python or other supported languages depending on the kernel) is written and executed. After execution, each code cell shows the results immediately below the cell. A notebook enables users to edit and execute each code cell independently, allowing them the flexibility to run cells in any order they prefer. This means that programmers are not obligated to execute their code sequentially from top to bottom or limited to making changes only at the end of the notebook. Instead, they can selectively modify and execute specific code cells at any point, fostering a more dynamic and interactive coding experience [21, 46].
Markdown Cell: Jupyter Notebook’s Markdown cells provide the narrative to a notebook. They use plain text with Markdown syntax to create formatted text and media, including headings, lists, links, images, and formatted text to explain and guide the reader through the notebook’s purpose, underlying logic, and methodology [115, 165]. This aspect of notebooks is crucial for making them comprehensive and user-friendly, especially when sharing with peers who may not be familiar with the code. Comments in traditional programming are quite similar to this Markdown cell. However, a Markdown cell generally contains more information than traditional textual comments [91, 100].
@@
Output Cell: The output of a code cell displays the results generated by the execution of the corresponding code cell. These can include textual data, tables, error messages, or visualizations created using libraries such as Matplotlib, [^4] https://pypi.org/project/matplotlib/ Seaborn, [^5] https://pypi.org/project/seaborn/ or Plotly. [^6] https://pypi.org/project/plotly/ The outputs are automatically updated whenever the associated code cell is executed. This capability makes Jupyter notebooks especially valuable for data science and research, as outputs can immediately visualize and communicate results alongside code, enhancing understanding and analysis [124, 118].
@@ 2.1.2 Execution Order
The execution order in a Jupyter notebook refers to the sequence in which the code cells are executed. Unlike traditional scripts, code cells in a notebook can be executed non-linearly, depending on the user’s needs and exploration path. The execution sequence is indicated by a number in brackets next to each code cell (e.g., In [1]:), which shows the most recent execution order (see Figure 1 for the execution number). Jupyter Notebook allows users to execute any code cell multiple times or out-of-order. This flexibility allows notebook users to test specific parts of their code, troubleshoot issues, or make adjustments without rerunning the entire notebook [140, 116, 55]. This type of execution in notebooks also helps the iterative nature of exploratory data analysis [33]. However, out-of-order execution can also cause problems, such as failure to load dependencies if they are not properly managed in the code [178]. Researchers found that about 36% of notebooks on GitHub did not execute in linear order [99].
@@ 2.2 Applications of Jupyter Notebooks
Jupyter Notebook is a popular tool for data analysis and data science, often used for interactive exploration during experiments [112, 67, 190]. It is considered an effective tool for both experienced and novice data scientists [36, 18]. Its flexibility in code execution facilitates short feedback cycles, which are essential for the iterative data analysis process [33]. While notebooks have been used with static visualizations, interactive visualizations can also be embedded and supported, as well as advanced visual analysis [106, 184]. These allow users to quickly try out different data analysis options and observe the results with minimal effort. Notebooks are also used for generating personalized data narrations over a given dataset for interactive data exploration [19] and sharing data science work through presentation slides [202, 176, 81, 107].
@@
Jupyter notebooks are widely used in the education sector and in academic teaching [6, 157, 113, 17]. They provide an intuitive and user-friendly platform for teaching programming concepts and designing course structures [129, 65]. Educators use Jupyter notebooks to create coding tutorials, assignments, and lecture materials that foster active learning and engagement among students [6, 4]. Research has shown that the use of Jupyter notebooks in educational settings significantly enhances the student’s understanding of course materials [9]. Additionally, studies demonstrate that Jupyter notebooks are particularly convenient for teaching data science, as they eliminate the need for students to install software locally or use specific machines [163]. Despite their popularity for managing coding-based assignments, automatic assignment grading in Jupyter notebooks presents challenges. The standard automated assessment grading tool for Jupyter notebooks, such as nbgrader [52]) has limitations; for example, it cannot automatically submit scores to a Learning Management System (LMS) like Canvas and does not provide automated feedback to students [95]. To address the automatic score submission, Malone et al. developed an automated assignment grading system for Jupyter notebooks that focuses on gamified cybersecurity exercises [94].
@@
Jupyter Notebook is popular for modeling and analyzing scientific tasks and performing experimental simulations [11, 142, 110, 164]. For example, Tran et al. developed a library for Jupyter Notebook that can set up and control additive manufacturing machines such as 3D printers [161]. Wilsdorf et al. presented a Jupyter Notebook extension that lends support to modelers by automatically specifying and running suitable simulation experiments [192]. Jupyter Notebook is also popular for analyzing GIS data to deal with big geospatial data [200, 199, 51]. For example, Valentine et al. implemented a data discovery studio for geoscience data discovery and exploration using Jupyter Notebook [162]. Yin et al. presented a cyberGIS framework to achieve data-intensive and scalable geospatial analytics using the Jupyter Notebook [200]. Jupyter notebooks are also used as a toolbox for interactive surface water mapping [108], geospatial environmental data processing [158], and astrophysical data-proximate analysis [62].
@@
Notebooks are often used in the healthcare sector to analyze complex medical data [48, 5, 8]. For example, Hao et al. from IBM Research developed customized healthcare data analysis pipelines in Jupyter Notebook to support healthcare users on user-friendly and reusable data analytics [53]. Almugbel et al. developed a tool that allows users to easily distribute their biomedical data analysis through notebooks uploaded to a GitHub repository or a private server [5]. The study provided four different Jupyter notebooks to infer differential gene expression, analyze cross-platform datasets, and process RNA sequence data. Biomedical researchers prefer notebooks to document and share their research with their community [5, 8]. For example, researchers used Jupyter Notebook as a co-design tool that combines static illustrations and interactive ML model explanations to predict health risk in diabetes care [8]. Furthermore, Llaunet et al. developed and shared a solution with centralized Jupyter Notebook code to support various medical applications such as medical image processing [79]. Beyond the above sectors, people use Jupyter Notebook in different areas. For example, journalists and media organizations use notebooks to analyze journalism content and visualize trends and patterns of newspaper data [148, 153].
@@ 3 Research Methodology
For our systematic literature review of software engineering research on Jupyter notebooks, we followed the guidelines proposed by Kitchenham [71]. The planning stage of our work includes two steps: (1) identifying the need for a systematic review and (2) developing the review protocol. In the conducting stage, based on the review protocol from the planning stage, we searched and selected the primary studies. We considered DBLP [^7] https://dblp.org/ as the search space, as it covers all the major software engineering journals and conference proceedings published by renowned publishers, e.g., IEEE, ACM, Springer, and Elsevier. Then, we extracted data from DBLP and synthesized the data. We selected 146 papers as our primary studies after completing all the steps in this process. Finally, in the reporting stage, we concluded the systematic review by reporting the collected data and findings. Figure 2 summarizes the steps of our methodology. In this section, we explain each step in more detail.

### v2

@@ Abstract
Method: We selected 199 relevant publications up to September 2025, following established systematic literature review guidelines. We explored publication trends, categorized them based on software engineering topics, and reported findings based on those topics.
Results: The most popular venues for publishing software engineering research on Jupyter notebooks are related to human-computer interaction instead of traditional software engineering venues. Researchers have addressed a wide range of software engineering topics on notebooks, such as code reuse, readability, and execution environment. Although reusability is one of the research topics for Jupyter notebooks, only 82 of the 199 studies can be reused based on their provided URLs. Additionally, most replication packages are not hosted on permanent repositories for long-term availability and adherence to open science principles.
@@
{highlights}
This research provides the first comprehensive systematic literature review on software engineering research specifically targeting Jupyter notebooks, identifying 199 primary studies published up to September 2025 and categorizing them into 11 core software engineering topics.
This research reveals that a large portion of the studies have been published outside traditional software engineering venues, with Human-Computer Interaction conferences like ACM Conference on Human Factors in Computing Systems (CHI) being the top publishing venues, highlighting the interdisciplinary nature of Jupyter Notebook research.
This research identifies a reusability gap in existing research, showing that only 82 out of 199 studies offer usable replication packages, and most are hosted on GitHub instead of permanent repositories, which violates open science best practices.
This research identifies that notebook-specific solutions for software engineering issues such as testing, refactoring, and documentation are relatively underexplored. Future directions include resolving duplicated execution numbers, refactoring inter-notebook clones, and generating grouped documentation for coherent-code cells are future directions derived from our study.
This research proposes the integration of modern AI-based solutions into Jupyter notebooks to support various software engineering topics, including code search and code generation. Additionally, future research should leverage advanced AI techniques (e.g., large language models), to improve conversational AI-powered assistants for automated code generation by multi-step workflow automation in data science notebooks.
Although the paper exceeds the recommended length due to the inclusion of detailed tables, figures, and categorized analyses (covering 11 topics and 21 subtopics), we believe that this extended content is essential for clearly and completely reporting our findings. As the first systematic literature review in this domain, we have carefully structured the paper to ensure readability. We believe the length is justified by the value and breadth of this paper’s contributions.
@@ 1 Introduction
In recent years, Jupyter Notebook has emerged as a powerful and versatile computing platform for data science, scientific research, and software development [perkel2018jupyter, rule2018exploration]. Its user-friendly interface has revolutionized the way researchers, developers, and data scientists conduct and communicate their work. Jupyter Notebook provides a unique platform for integrating code, visualizations, and narrative text, fostering reproducibility, and facilitating collaborative exploration of data and algorithms [wang2019data]. Seamlessly incorporating code, data, and output into a single file makes the Jupyter Notebook ideal for data analysis, scientific computing, and machine learning (ML) tasks. Moreover, its support for multiple programming languages, including Python, R, and Julia, makes it versatile and widely accessible to a diverse community of researchers, data scientists, and educators. However, despite their popularity, notebooks have been associated with several challenges, such as low reproducibility rates, problems with their execution environments, and excessive code duplication [grotov2022large].
As Jupyter notebooks are often created and maintained by users without a software engineering background [simmons2020large], there has been a growing interest from software engineering researchers to help notebook users integrate best software engineering practices into their notebooks. As a result, there has been a growing body of software engineering research that targets Jupyter Notebook. In this paper, we conduct a systematic literature review (SLR) of such research and the problems addressed by it.
We systematically searched academic publications to identify relevant studies related to software engineering research on Jupyter Notebook. We followed Kitchenham’s guidelines [kitchenham2004procedures] for our SLR to ensure rigor and impartiality. We comprehensively searched all the papers on Jupyter Notebook indexed in the ACM digital library, IEEE Xplore, Elsevier, Springer Link, and DBLP databases published until September 2025. Then, we filtered the papers focused on software engineering research on Jupyter notebooks. The review process involved several steps, including screening the titles and abstracts of the identified studies, assessing their relevance based on the inclusion criteria, and extracting relevant data from the selected studies. We finally selected 199 papers as our primary studies. The overall characterization of our review based on Cooper’s taxonomy [Cooper1988Taxonomy] is shown in Table 1. During our SLR, we focused on the following two research questions (RQs):
Table 1: Characterization of this SLR based on Cooper’s taxonomy [Cooper1988Taxonomy].
Dimension Our SLR
Focus Research outcomes, research methods, practices or applications
Goal Integration, identification of central issues
Perspective Neutral representation
Coverage Exhaustive review with selective citation (“Jupyter” or “notebook” in the title) and limited to software engineering research
Organization Conceptual (topic-centric; see Table 5)
Audience Specialized scholars in software engineering, practitioners
@@
    RQ1: How much software engineering research on Jupyter Notebook has been published? We analyzed current trends in the publication of software engineering research on notebooks to help future researchers identify potential venues for their work. The number of publications has gradually increased over the years. Most (104 of 199 studies) software engineering research on notebooks has been published at conferences, indicating a fast-moving field. Our research found that other venues beyond core software engineering conferences published the highest number of studies; for example, the ACM Conference on Human Factors in Computing Systems (CHI) published 20 studies. Conversely, the International Conference on Software Engineering (ICSE) published seven studies on software engineering research on notebooks. According to our study, a significant number of the articles we examined, specifically 37 out of 199 (18.8%), had affiliations with industry institutions such as Microsoft Research (13 studies) or IBM Research Lab (6 studies). This highlights the practical importance and real-world applications of the research findings. We identified 87 studies that provided URLs to replication packages. Of these, 82 URLs were deemed reusable, while 5 were not, due to missing source code or unclear execution instructions. Most replication packages (69 studies) are hosted in GitHub repositories, which is against the best practices for open science in software engineering [Mendez_2020] as such repositories can disappear over time. Only 17 replication packages were hosted in a permanent repository, such as Zenodo or Figshare.
@@
    RQ2: Which software engineering topics are being studied in software engineering research on Jupyter Notebook? We categorized the primary studies into 11 software engineering topics: code reuse and provenance, managing the computational environment and workflow, readability of notebooks, datasets of notebooks, documentation of notebooks, testing and debugging, visualization in notebooks, best practices, cell execution order, notebook code generation, and supporting other programming paradigms. Our findings indicate that managing computational environment and workflow is the most extensively researched topic related to Jupyter Notebook, with 42 studies focusing on it. This is followed by code reuse and provenance with 41 studies, and documentation of notebooks with 22 studies. These results suggest that researchers are primarily concerned with code cells in notebooks, which often require significant human effort to understand. We also identified 22 publicly available Jupyter Notebook datasets used in the studies, most of which were sourced from Kaggle and GitHub repositories.
@@ 2.1.1 Cell-Based Structure
Code Cell: Code cells are the fundamental components of a Jupyter Notebook, where the source code (Python or other supported languages depending on the kernel) is written and executed. After execution, each code cell shows the results immediately below the cell. A notebook enables users to edit and execute each code cell independently, allowing them the flexibility to run cells in any order they prefer. This means that programmers are not obligated to execute their code sequentially from top to bottom or limited to making changes only at the end of the notebook. Instead, they can selectively modify and execute specific code cells at any point, fostering a more dynamic and interactive coding experience [chattopadhyay2020whats, grotov2022large].
Markdown Cell: Jupyter Notebook’s Markdown cells provide the narrative to a notebook. They use plain text with Markdown syntax to create formatted text and media, including headings, lists, links, images, and formatted text to explain and guide the reader through the notebook’s purpose, underlying logic, and methodology [pimentel2019large, venkatesh2021automated]. This aspect of notebooks is crucial for making them comprehensive and user-friendly, especially when sharing with peers who may not be familiar with the code. Comments in traditional programming are quite similar to this Markdown cell. However, a Markdown cell generally contains more information than traditional textual comments [liu2021haconvgnn, mondal2023cell2doc].
@@
Output Cell: The output of a code cell displays the results generated by the execution of the corresponding code cell. These can include textual data, tables, error messages, or visualizations created using libraries such as Matplotlib, [^4] https://pypi.org/project/matplotlib/ Seaborn, [^5] https://pypi.org/project/seaborn/ or Plotly. [^6] https://pypi.org/project/plotly/ The outputs are automatically updated whenever the associated code cell is executed. This capability makes Jupyter notebooks especially valuable for data science and research, as outputs can immediately visualize and communicate results alongside code, enhancing understanding and analysis [raghunandan2023measuring, psallidas2022data].
@@ 2.1.2 Execution Order
The execution order in a Jupyter notebook refers to the sequence in which the code cells are executed. Unlike traditional scripts, code cells in a notebook can be executed non-linearly, depending on the user’s needs and exploration path. The execution sequence is indicated by a number in brackets next to each code cell (e.g., In [1]:), which shows the most recent execution order (see Figure 1 for the execution number). Jupyter Notebook allows users to execute any code cell multiple times or out-of-order. This flexibility allows notebook users to test specific parts of their code, troubleshoot issues, or make adjustments without rerunning the entire notebook [sato2022comparing, pimentel2021understanding, head2019managing]. This type of execution in notebooks also helps the iterative nature of exploratory data analysis [fangohr2019data]. However, out-of-order execution can also cause problems, such as failure to load dependencies if they are not properly managed in the code [wang2020restoring]. Researchers found that about 36% of notebooks on GitHub did not execute in linear order [michael2022keeping].
@@ 2.2 Applications of Jupyter Notebooks
Jupyter Notebook is a popular tool for data analysis and data science, often used for interactive exploration during experiments [perkel2018jupyter, kery2018story, werner2021bridging]. It is considered an effective tool for both experienced and novice data scientists [fruchart2022jupyter, chandel2022training]. Its flexibility in code execution facilitates short feedback cycles, which are essential for the iterative data analysis process [fangohr2019data]. While notebooks have been used with static visualizations, interactive visualizations can also be embedded and supported, as well as advanced visual analysis [ono2021interactive, wang2023supernova]. These allow users to quickly try out different data analysis options and observe the results with minimal effort. Notebooks are also used for generating personalized data narrations over a given dataset for interactive data exploration [chanson2022generating] and sharing data science work through presentation slides [zheng2022telling, wang2023slide4n, li2023notable, ouyang2024noteplayer].
@@
Jupyter notebooks are widely used in the education sector and in academic teaching [amoudi2023interactive, tan2021nascent, petersohn2023kopplung, casseau2023moon]. They provide an intuitive and user-friendly platform for teaching programming concepts and designing course structures [reades2020teaching, kastner2020teaching]. Educators use Jupyter notebooks to create coding tutorials, assignments, and lecture materials that foster active learning and engagement among students [amoudi2023interactive, al2022jupyter]. Research has shown that the use of Jupyter notebooks in educational settings significantly enhances the student’s understanding of course materials [barba2019teaching]. Additionally, studies demonstrate that Jupyter notebooks are particularly convenient for teaching data science, as they eliminate the need for students to install software locally or use specific machines [van2020jupyter]. Despite their popularity for managing coding-based assignments, automatic assignment grading in Jupyter notebooks presents challenges. The standard automated assessment grading tool for Jupyter notebooks, such as nbgrader [hamrick2016creating]) has limitations; for example, it cannot automatically submit scores to a Learning Management System (LMS) like Canvas and does not provide automated feedback to students [manzoor2020auto]. To address the automatic score submission, Malone et al. developed an automated assignment grading system for Jupyter notebooks that focuses on gamified cybersecurity exercises [malone2023securely].
@@
Jupyter Notebook is popular for modeling and analyzing scientific tasks and performing experimental simulations [beg2021using, savira2021writing, penuela2021open, vandewalle2019integrating]. For example, Tran et al. developed a library for Jupyter Notebook that can set up and control additive manufacturing machines such as 3D printers [tran2023imprimer]. Wilsdorf et al. presented a Jupyter Notebook extension that lends support to modelers by automatically specifying and running suitable simulation experiments [wilsdorf2023nbsimgen]. Jupyter Notebook is also popular for analyzing GIS data to deal with big geospatial data [yin2017cybergis, yin2019cybergis, haedrich2023integrating]. For example, Valentine et al. implemented a data discovery studio for geoscience data discovery and exploration using Jupyter Notebook [valentine2021earthcube]. Yin et al. presented a cyberGIS framework to achieve data-intensive and scalable geospatial analytics using the Jupyter Notebook [yin2017cybergis]. Jupyter notebooks are also used as a toolbox for interactive surface water mapping [owusu2022pygee], geospatial environmental data processing [terlych2021jupyter], and astrophysical data-proximate analysis [juneau2021jupyter].
@@
Notebooks are often used in the healthcare sector to analyze complex medical data [gruning2017jupyter, almugbel2018reproducible, ayobi2023computational]. For example, Hao et al. from IBM Research developed customized healthcare data analysis pipelines in Jupyter Notebook to support healthcare users on user-friendly and reusable data analytics [hao2017developing]. Almugbel et al. developed a tool that allows users to easily distribute their biomedical data analysis through notebooks uploaded to a GitHub repository or a private server [almugbel2018reproducible]. The study provided four different Jupyter notebooks to infer differential gene expression, analyze cross-platform datasets, and process RNA sequence data. Biomedical researchers prefer notebooks to document and share their research with their community [almugbel2018reproducible, ayobi2023computational]. For example, researchers used Jupyter Notebook as a co-design tool that combines static illustrations and interactive ML model explanations to predict health risk in diabetes care [ayobi2023computational]. Furthermore, Llaunet et al. developed and shared a solution with centralized Jupyter Notebook code to support various medical applications such as medical image processing [launet2023federating]. Beyond the above sectors, people use Jupyter Notebook in different areas. For example, journalists and media organizations use notebooks to analyze journalism content and visualize trends and patterns of newspaper data [showkat2021stories, stark2016towards].
@@ 3 Research Methodology
For our systematic literature review of software engineering research on Jupyter notebooks, we followed the guidelines proposed by Kitchenham [kitchenham2004procedures]. The planning stage of our work includes two steps: (1) identifying the need for a systematic review and (2) developing the review protocol. In the conducting stage, based on the review protocol from the planning stage, we searched and selected the primary studies. We considered five different digital libraries (IEEE Xplore, [^7] https://ieeexplore.ieee.org/ ACM Digital Library, [^8] https://dl.acm.org/ Elsevier ScienceDirect, [^9] https://www.sciencedirect.com/ Springer Nature Link [^10] https://link.springer.com/ and DBLP [^11] https://dblp.org/) as the search space, as those cover all the major software engineering journals and conference proceedings published by renowned publishers. Then, we extracted and synthesized the data. We selected 199 papers as our primary studies after completing all the steps in this process. Finally, in the reporting stage, we concluded the systematic review by reporting the collected data and findings. Figure 2 summarizes the steps of our methodology. In this section, we explain each step in more detail.